OpenStudy (anonymous):

Need help. Don't know how to do this. Solve for D F=g(mM/d^2)

OpenStudy (anonymous):

also, I apologize, my laptop is really slow

OpenStudy (anonymous):

I tried solving myself but the answer was wrong I'm completely lost.

OpenStudy (radar):

I don't see a D, do you mean the d

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the little d

OpenStudy (anonymous):

F= g(mM/d^2) \[d^2=\frac{g m M}{F} \]\[d=\sqrt{\frac{g m M}{F}} \]

OpenStudy (radar):

There it is, robtobey to the rescue.

OpenStudy (anonymous):

oh wow really that simple?